In 200cm3 of water are added 10 cm3 of KOH and the solvate formed shows pH = 12.2. How much was the initial concentration of KOH?
1
Expert's answer
2019-04-29T03:17:04-0400

"c(KOH)=c(OH-)=10^{pH-14}=10^{-1.8}=0.01585 mol\/L"

"n(KOH)=c*V= 0.01585 mol\/L*0.21L =0.0033mol"

"c(KOH)= n\/V = 0.0033mol \/ 0.01L=0.33mol\/L"